OpenStudy (anonymous):

The length of a rectangle is 20 in and the width is 35 in. The rectangle is then dilated by a scale factor of 1/3. What are the new dimensions?

OpenStudy (anonymous):

We are basically increasing each dimension by 1/3. Let's call the new dimenions x and y, i.e x will be the new dilated with and y will be the new dilated length. \[\bf x = 35 +35\left( \frac{ 1 }{ 3 } \right)\]Similarly:\[\bf y=20+20\left( \frac{ 1 }{ 3 } \right)\]Can you find the length and width?

OpenStudy (anonymous):

don't you just solve them and then you find the length and width ?

OpenStudy (anonymous):

x = 46.7 and y would = 26.7 ?